# On a pinning model in correlated Gaussian random environments

### v1

@@ Abstract.
We consider a pinning model in correlated Gaussian random environments. For the model that is disorder relevant, we study its intermediate disorder regime and show that the rescaled partition functions converge to a non-trivial continuum limit in the Skorohod setting and in the Stratonovich setting, respectively. Our results partially confirm the prediction of Weinrib and Halperin [21] for disorder relevance/irrelevance.
@@ 1. Introduction
In this paper, we study a pinning model in Gaussian random environments (also called the disorder), of which the correlation has a power-law decay. We shall show that, under proper scalings, the rescaled partition functions converge weakly to a continuum limit which can be represented by a chaos expansion. Our results complement those of Berger [2] and are consistent with the prediction of Weinrib and Halperin [21] on the disorder relevance/irrelevance for physical systems with correlated disorder.
@@ 1.1. The disordered pinning model
The disordered pinning model is a type of random polymer models, which is defined via a Gibbs transform of the distribution of a renewal process. Pinning models are well suited models for studying phase transitions and critical phenomena and also have received widespread attention in physics, chemistry and biology. We refer to the monographs of Giacomin [9] and [10] for more details.
@@
where $\alpha>0$ and $L(\cdot):\mathbb{R}^{+}\to\mathbb{R}^{+}$ is a slowly varying function, i.e., $\lim_{t\to\infty}L(at)/L(t)=1$ for any $a>0$ (see [5] for more details about slowly varying functions). The expectation with respect to ${\mathbb{P}}_{\tau}$ is denoted by $\mathbb{E}_{\tau}$. Note that $\mathbb{E}_{\tau}[\tau_{1}]<\infty$ for $\alpha>1$ and $\mathbb{E}_{\tau}[\tau_{1}]=\infty$ for $\alpha<1$, and by [5, Proposition 1.5.9a], for $\alpha=1$,
@@
is also slowly varying (whether it is finite or not depending on $L(\cdot)$). Hence, the renewal process $\tau$ can be classified by the following conditions:
(1) $10<{\alpha}<$ 1; (2) $\mathbb{E}_{\tau}[\tau_{1}]<\infty$; (3) $\alpha=1$ and $\mathbb{E}_{\tau}[\tau_{1}]=\infty$. (1.3)
Noting that our renewal process $\tau$ is aperiodic since $L(n)>0$ for all $n$, we have the renewal theorem (see [8] and [5, Theorem 8.7.5]):
${\mathbb{P}}_{\tau}(n\in\tau)\sim\begin{cases}\dfrac{C_{\alpha}}{L(n)n^{1-\alpha}},&\text{if }0<\alpha<1~{}(C_{\alpha}=\frac{\alpha\sin(\pi\alpha)}{\pi}),\\[10.0pt] \dfrac{1}{\mathbb{E}[\tau_{1}]},&\text{if }\mathbb{E}[\tau_{1}]<\infty,\\[10.0pt] \dfrac{1}{l(n)},&\text{if }\alpha=1\text{ and }\mathbb{E}[\tau_{1}]=\infty,\end{cases}\quad\qquad\text{as}~{}n\to\infty.$ (1.4)
Let $\omega=\{\omega_{n}\text{,}~{}n\in\mathbb{N}\}$ be a family of Gaussian random variables independent of $\tau$, representing the random environments. The probability and expectation for $\omega$ are denoted by ${\mathbb{P}}_{\omega}$ and $\mathbb{E}_{\omega}$, respectively. We assume that
$\mathbb{E}_{\omega}[\omega_{n}]=0\quad\text{and}\quad\mathbb{E}_{\omega}[\omega_{n}\omega_{n^{\prime}}]=\gamma(n-n^{\prime}),\quad\text{for }n,n^{\prime}\in\mathbb{N},$ (1.5)
@@
Denote $\Omega_{N}:=\Big{\{}\frac{1}{N}\text{,}\frac{2}{N}\text{,}\cdots\text{,}\frac{N-1}{N}\text{,}1\Big{\}}$ and $\delta_{x}^{(N)}=\delta_{x}:=\mathbf{1}_{\{Nx\in\tau\}}$. The disordered pinning model with free boundary is defined by
$\frac{{\mathrm{d}}{\mathbb{P}}_{N,\beta}^{\omega,h}}{{\mathrm{d}}{\mathbb{P}}_{\tau}}(\tau):=\frac{1}{Z_{N,\beta}^{\omega,h}}\exp\Big{\{}\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}\Big{\}},$
@@
1.2. Disorder relevance/irrelevance and the Weinrib-Halperin prediction
@@
When the disorder is i.i.d., the Harris criterion [11] indicates that disorder is irrelevant if ${\alpha}<\frac{1}{2}$ and is relevant if ${\alpha}>\frac{1}{2}$ (with $\alpha=\frac{1}{2}$ being the marginal case). When the disorder is correlated and the correlation is of the form $n^{-p}$ ($p>0$ and $n$ is the distance between two points), Weinrib and Halperin [21] predicted that the disorder should be relevant if ${\alpha}>\min(p,1)/2$ and irrelevant if ${\alpha}<\min(p,1)/2$. Berger [2] confirmed the Weinrib-Halperin prediction for the case $p>1$ (note that in this case the Harris criterion still applies), and showed that the case $p<1$ is different in the sense that classical methods (see, e.g., [10]) in the study of disorder relevance/irrelevance by comparing critical points/exponents no longer apply: when $p<1$, the disordered system does not possess a localization/delocalization phase transition, since the annealed free energy is always infinite and the quenched free energy is always strictly positive for all $h$ as long as $\beta>0$, and thus the system is always localized (see Theorem 2.5 and Section 2.3 in [2]). This phenomenon is due to the fact, as pointed out by Berger [3], that when $p<1$, there are large and frequent regions in the environments with arbitrarily high attractions (called the “infinite disorder” in [3]). These results hint at scalings of $\beta$ and $h$ when studying disorder relevance/irrelevance for the case $p<1$.
@@
On the other hand, Caravenna, Sun and Zygouras [6] proposed a new perspective on disorder relevance/irrelevance: noting that disorder relevance means that an arbitrarily small amount of disorder (i.e., for any $\beta>0$) changes the features of the underlying homogeneous model, it is possible that if the strength of disorder is tuned down to zero at a proper rate (i.e., $\beta_{N}\to 0$) as the system size tends to infinity, then the disorder will persist in the resulting continuum model. This phenomenon has been shown for the i.i.d. disorder by studying weak limits of rescaled partition functions for the directed polymer model in [1] and for pinning models, directed polymer models and the Ising model in a unified way in [6]. For related results on directed polymers, we refer to [18, 7] for the case when the disorder is correlated in space but independent in time, to [19] for the case when the disorder is correlated in time but independent in space, and to [20] for the case when the disorder is space-time correlated. In those models, the Weinrib-Halperin prediction is not clear (the decay rates in space and in time of the correlation of the disorder in [20] do not agree).
@@
Inspired by the above-mentioned works, we study the weak convergence of the rescaled partition functions. In our setting, noting that $p=2-2H\in(0,1)$ with $H\in(\frac{1}{2},1)$, the Weinrib-Halperin prediction suggests that the disorder relevent if $\alpha>1-H$, which coincides with the condition for the existence of the weak limit for each chaos (see Proposition 3.1).
@@
Finally, we make a few more remarks on disorder relevance/irrelevance for pinning models in correlated random environment. Poisat [17] showed that the Harris criterion still applies to the pinning model in random environments with finite range correlation. When disorders have long range correlation, Berger and Poisat [4] conformed the Weinrib-Halperin prediction for the case when the correlations are summable, which extended the result in [2].
@@ 1.3. Main results
Noting that
@@
we first consider the following Wick ordered partition function
$\tilde{Z}_{N,\beta}^{\omega,h}:=\mathbb{E}_{\tau}\Big{[}\exp\Big{\{}\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}-\frac{1}{2}\mathrm{Var}_{\omega}\Big{(}\sum_{x\in\Omega_{N}}\beta\omega_{Nx}\delta_{x}\Big{)}\Big{\}}\Big{]}.$ (1.10)
Note that the term $\frac{1}{2}\mathrm{Var}_{\omega}(\sum_{x\in\Omega_{N}}\beta\omega_{Nx}\delta_{x})=\frac{1}{2}\beta^{2}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n,m\in\tau\}}$ is a direct normalization of the partition function, instead of the Gibbs measure itself, since it depends on the realization of $\tau$. We adopt this normalization because the disorders are correlated, and thus a polynomial chaos expansion as has been done in [1, 6] cannot be employed for our pinning model. The normalization in the Wick ordered partition function (1.10) results in a Wick exponential, which allows us to perform a Wick expansion (see Section 3 and also the discussion in [13]). We point out that the Wick ordered partition function (1.10) reduces to [6, equation (3.3)] when the disorders are i.i.d., noting that $\frac{1}{2}\mathrm{Var}_{\omega}(\sum_{x\in\Omega_{N}}\beta\omega_{Nx}\delta_{x})=\frac{\beta^{2}}{2}\sum_{n=1}^{N}\mathbf{1}_{\{n\in\tau\}}$.
@@
$\beta_{N}=\begin{cases}\hat{\beta}\dfrac{L(N)}{N^{\alpha+H-1}},&\text{if }0<\alpha<1,\\[10.0pt] \hat{\beta}\dfrac{1}{N^{H}},&\text{if }\mathbb{E}[\tau_{1}]<\infty,\\[10.0pt] \hat{\beta}\dfrac{l(N)}{N^{H}},&\text{if }\alpha=1\text{ and }\mathbb{E}[\tau_{1}]=\infty,\end{cases}h_{N}=\begin{cases}\hat{h}\dfrac{L(N)}{N^{\alpha}},&\text{if }0<\alpha<1,\\[10.0pt] \hat{h}\dfrac{1}{N},&\text{if }\mathbb{E}[\tau_{1}]<\infty,\\[10.0pt] \hat{h}\dfrac{l(N)}{N},&\text{if }\alpha=1\text{ and }\mathbb{E}[\tau_{1}]=\infty,\end{cases}$ (1.11)
@@
Let $\phi_{k}(t_{1},\dots,t_{k})$ be a symmetric function, defined for $10<t_{1}<\cdots<t_{k}<$ 1 by
$\phi_{k}(t_{1},\dots,t_{k}):=\begin{cases}\displaystyle\frac{C_{\alpha}^{k}}{t_{1}^{1-\alpha}(t_{2}-t_{1})^{1-\alpha}\cdots(t_{k}-t_{k-1})^{1-\alpha}},&\text{ if }0<\alpha<1,\\[10.0pt] \displaystyle\Big{(}\frac{1}{\mathbb{E}[\tau_{1}]}\Big{)}^{k},&\text{ if }\mathbb{E}[\tau_{1}]<\infty,\\[10.0pt] \displaystyle 1,&\text{ if }\alpha=1\text{ and }\mathbb{E}[\tau_{1}]=\infty,\end{cases}$ (1.12)
with $C_{\alpha}=\frac{\alpha\sin(\pi\alpha)}{\pi}$ as in (1.4).
Now we present our main results for the Wick ordered partition function $\tilde{Z}_{N,\beta}^{\omega,h}$ given in (1.10) and for the original partition function $Z_{N,\beta}^{\omega,h}$ given in (1.8), respectively.
@@
Let the renewal process $\tau$ and the disorder $\omega$ be given as in Section 1.1. Let H and $\alpha$ be parameters satisfying
$H\in\Big{(}\frac{1}{2},1\Big{)}\text{ and }~{}\alpha>\frac{1}{2}.$
Recalling the Wick ordered partition function $\tilde{Z}^{\omega,h}_{N,\beta}$ given in (1.10) and the scalings (1.11), we have that
@@
where $\tilde{Z}_{\hat{\beta},\hat{h}}$ is given by the $L^{2}$-convergent series:
where $\phi_{k}$ is given by (1.12), $\dot{W}$ is Gaussian noise with $\mathbb{E}[\dot{W}(t)\dot{W}(s)]=|t-s|^{2H-2}$ and the stochastic integral is of Skorohod type.
@@
Let the renewal process $\tau$ and the disorder $\omega$ be given as in Section 1.1. Let H and $\alpha$ be parameters satisfying
$H\in\Big{(}\frac{1}{2},1\Big{)}\text{ and }~{}\alpha+H>\frac{3}{2}.$
Recalling the partition function $Z^{\omega,h}_{N,\beta}$ given in (1.8) and the scalings (1.11), we have that
$Z^{\omega,h_{N}}_{N,\beta_{N}}\overset{(d)}{\longrightarrow}Z_{\hat{\beta},\hat{h}}~{},~{}as~{}N\to\infty,$
where $Z_{\hat{\beta},\hat{h}}$ is given by the $L^{2}$-convergent series:
@@
The results in Theorems 1.1 and 1.2 also hold for the conditioned partition functions
$\tilde{Z}_{N}^{\text{c}}=\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N},\text{c}}:=\mathbb{E}_{\tau}\Big{[}\exp\Big{\{}\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}-\frac{1}{2}\mathrm{Var}\Big{[}\sum_{x\in\Omega_{N}}(\beta\omega_{Nx})\delta_{x}\Big{]}\Big{\}}\Big{|}N\in\tau\Big{]}$
and
@@
with the integrand $\phi_{k}$ in (1.13) and (1.14) being replaced by a symmetric function $\phi_{k}^{c}$, defined for $10<t_{1}<\cdots<t_{k}<$ 1 by
$\phi_{k}^{c}(t_{1},\dots,t_{k}):=\begin{cases}\displaystyle\frac{C_{\alpha}^{k}}{t_{1}^{1-\alpha}(t_{2}-t_{1})^{1-\alpha}\cdots(t_{k}-t_{k-1})^{1-\alpha}(1-t_{k})^{1-\alpha}},&\text{ if }1-H<\alpha<1,\\[10.0pt] \displaystyle\Big{(}\frac{1}{\mathbb{E}[\tau_{1}]}\Big{)}^{k},&\text{ if }\mathbb{E}[\tau_{1}]<\infty,\\[10.0pt] \displaystyle 1,&\text{ if }\alpha=1\text{ and }\mathbb{E}[\tau_{1}]=\infty.\end{cases}$

### v2

@@ Abstract.
We consider a pinning model in correlated Gaussian random environments. For the model that is disorder relevant, we study its intermediate disorder regime and show that the rescaled partition functions converge to a non-trivial continuum limit in the Skorohod setting and in the Stratonovich setting, respectively. Our results partially confirm the Weinrib–Halperin prediction for disorder relevance/irrelevance.
@@ 1. Introduction
In this paper, we study a pinning model in Gaussian random environments (also called the disorder), of which the correlation has a power-law decay. We shall show that, under proper scalings, the rescaled partition functions converge weakly to continuum limits which can be represented by chaos expansions. Our results complement those of [4] and are consistent with the prediction of [33] on the disorder relevance/irrelevance for physical systems with correlated disorders.
@@ 1.1. The disordered pinning model
The disordered pinning model is a type of random polymer models, which is defined via a Gibbs transform of the distribution of a renewal process. Pinning models are well suited models for studying phase transitions and critical phenomena and also have received widespread attention in physics, chemistry and biology. We refer to the monographs of [15] and [16] for more details.
@@
where $\alpha>0$ and $L(\cdot):\mathbb{R}^{+}\to\mathbb{R}^{+}$ is a slowly varying function, i.e., $\lim_{t\to\infty}L(at)/L(t)=1$ for any $a>0$ (see [7] for more details about slowly varying functions). The expectation with respect to ${\mathbb{P}}_{\tau}$ is denoted by $\mathbb{E}_{\tau}$. Note that $\mathbb{E}_{\tau}[\tau_{1}]<\infty$ for $\alpha>1$ and $\mathbb{E}_{\tau}[\tau_{1}]=\infty$ for $\alpha<1$, and by [7, Proposition 1.5.9a], for $\alpha=1$,
@@
is also slowly varying (whether it is finite or not depends on $L(\cdot)$). Hence, the renewal process $\tau$ can be classified by the following conditions:
(1) $0<{\alpha}<1$; (2) $\mathbb{E}_{\tau}[\tau_{1}]<\infty$; (3) $\alpha=1$ and $\mathbb{E}_{\tau}[\tau_{1}]=\infty$. (1.3)
Noting that our renewal process $\tau$ is aperiodic since $L(n)>0$ for all $n$, we have the renewal theorem (see [14] and [7, Theorem 8.7.5]):
${\mathbb{P}}_{\tau}(n\in\tau)\sim\begin{cases}\dfrac{C_{\alpha}}{L(n)n^{1-\alpha}},&\text{if }0<\alpha<1~\Big(C_{\alpha}=\frac{\alpha\sin(\pi\alpha)}{\pi}\Big),\\[10.0pt] \dfrac{1}{\mathbb{E}[\tau_{1}]},&\text{if }\mathbb{E}[\tau_{1}]<\infty,\\[10.0pt] \dfrac{1}{l(n)},&\text{if }\alpha=1\text{ and }\mathbb{E}[\tau_{1}]=\infty,\end{cases}\quad\qquad\text{as}~n\to\infty.$ (1.4)
Let $\omega=\{\omega_{n}\text{,}~n\in\mathbb{N}\}$ be a family of Gaussian random variables independent of $\tau$, representing the random environment. The probability and expectation for $\omega$ are denoted by ${\mathbb{P}}_{\omega}$ and $\mathbb{E}_{\omega}$, respectively. We assume that
[…]
@@
Denote $\Omega_{N}:=\{\frac{1}{N}\text{,}\frac{2}{N}\text{,}\cdots\text{,}\frac{N-1}{N}\text{,}1\}$ and $\delta_{x}^{(N)}=\delta_{x}:=\mathbf{1}_{\{Nx\in\tau\}}$. The disordered pinning model with free boundary is defined via Gibbs transform by
$\frac{{\mathrm{d}}{\mathbb{P}}_{N,\beta}^{\omega,h}}{{\mathrm{d}}{\mathbb{P}}_{\tau}}(\tau):=\frac{1}{Z_{N,\beta}^{\omega,h}}\exp\Big\{\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}\Big\},$ (1.8)
@@
1.2. Disorder relevance/irrelevance and the Weinrib–Halperin prediction
When the disorder is i.i.d., the Harris criterion [17] indicates that disorder is irrelevant if ${\alpha}<\frac{1}{2}$ and is relevant if ${\alpha}>\frac{1}{2}$ (with $\alpha=\frac{1}{2}$ being the marginal case). Here, “disorder is relevant” means that for any $\beta>0$, the behavior of $\tau$ under ${\mathbb{P}}_{N,\beta}^{\omega,h}$ differs drastically from that under ${\mathbb{P}}_{\tau}$. To be specific, some crucial quantities, such as the correlation length and the critical point shift will be changed by introducing disorder in the Gibbs transform (1.8). When disorder is irrelevant, the behavior of $\tau$ under ${\mathbb{P}}_{N,\beta}^{\omega,h}$ is comparable to that under ${\mathbb{P}}_{\tau}$, at least for small enough $\beta>0$. Readers may refer to [9, 8, 2, 11] for more detailed discussion about disorder relevance and irrelevance.
@@ 1.2. Disorder relevance/irrelevance and the Weinrib–Halperin prediction
When the disorder is correlated and the correlation is of the form $n^{-p}$ ($p>0$ and $n$ is the distance between two points), [33] predicted that the disorder should be relevant if ${\alpha}>\min(p,1)/2$ and irrelevant if ${\alpha}<\min(p,1)/2$. [4] confirmed the Weinrib–Halperin prediction for the case $p>1$ (note that in this case the Harris criterion still applies), and showed that the case $p<1$ is different in the sense that classical methods (see, e.g., [16]) in the study of disorder relevance/irrelevance by comparing critical points/exponents no longer apply: when $p<1$, the disordered system does not possess a localization/delocalization phase transition, since the annealed free energy is always infinite and the quenched free energy is always strictly positive for all $h$ as long as $\beta>0$, and thus the system is always localized (see Theorem 2.5 and Section 2.3 in [4]). This phenomenon is due to the fact, as pointed out by [5], that when $p<1$, there are large and frequent regions in the environments with arbitrarily high attractions (called the “infinite disorder” in [5]). These results indicate that different methods are needed to study disorder relevance/irrelevance for the case $p<1$.
@@
On the other hand, [9] proposed a new perspective on disorder relevance/irrelevance: it is possible that if the strength of disorder is tuned down to zero at a proper rate (i.e., $\beta_{N}\to 0$) as the system size tends to infinity, then the disorder will persist in the resulting continuum model. This phenomenon has been shown for the i.i.d. disorder by studying weak limits of rescaled partition functions for the directed polymer model in [1] and for pinning models, directed polymer models and the Ising model in a unified way in [9]. For related results on directed polymers, we refer to [30, 13] for the case when the disorder is correlated in space but independent in time, to [29] for the case when the disorder is correlated in time but independent in space, and to [31] for the case when the disorder is space-time correlated. In those models, the Weinrib–Halperin prediction is not clear (the decay rates in space and in time of the correlation of the disorder in [31] do not agree).
@@
Inspired by the above-mentioned works, we study the weak convergence of the rescaled partition functions. In our setting, noting that $p=2-2H\in(0,1)$ with $H\in(\frac{1}{2},1)$, the Weinrib–Halperin prediction suggests that disorder is relevant if $\alpha>1-H$, which coincides with the condition for the existence of the weak limit for each chaos (see Proposition 3.1). When $H\in(\frac{1}{2},1)$, under the stronger condition $\alpha\geq\frac{1}{2}$, we establish in Theorem 1.1 that the rescaled Wick-ordered partition functions converge weakly to an $L^{2}$ random variable. We further conjecture that for $H\in(\frac{1}{2},1)$, these rescaled Wick-ordered partition functions converge weakly to an $L^{1}$ random variable under the Weinrib–Halperin condition $\alpha>1-H$ (see Remark 1.6 and Section 3.2 for further details).
@@
Finally, we make a few more remarks on disorder relevance/irrelevance for pinning models in correlated random environment. [26] showed that the Harris criterion still applies to the pinning model in random environments with finite range correlation. When disorders have long range correlation, [3] confirmed the Weinrib–Halperin prediction for the case when the correlations are summable, which extended the result in [4], and [27] obtained sharp asymptotics of the annealed free energy for doubly summable or exponentially decaying correlations. We also mention that [12] studied a disordered pinning model, where the random environment is given by another independent renewal process. In that case, the random environment is correlated and it is different from the more typical setting where the random environment is a family of random variables. They found that in their model, the chaos expansion heuristics and the Weinrib–Halperin prediction provide different phase diagram.
@@ 1.3. Main results
Rather than studying the Gibbs measure directly, which is highly involved, we focus on its partition function, which already encapsulates sufficient information about the model. Noting that
@@
we first consider the following Wick-ordered partition function
$\tilde{Z}_{N,\beta}^{\omega,h}:=\mathbb{E}_{\tau}\Big[\exp\Big\{\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}-\frac{1}{2}\mathrm{Var}_{\omega}\Big(\sum_{x\in\Omega_{N}}\beta\omega_{Nx}\delta_{x}\Big)\Big\}\Big],$ (1.11)
where the term $\frac{1}{2}\mathrm{Var}_{\omega}(\sum_{x\in\Omega_{N}}\beta\omega_{Nx}\delta_{x})=\frac{1}{2}\beta^{2}\sum_{n,m=1}^{N}\gamma(n-m)\mathbf{1}_{\{n,m\in\tau\}}$ is a direct normalization of the partition function (we call it Wick normalization), which depends on the realization of $\tau$. We adopt this normalization because disorder is correlated, and thus a polynomial chaos expansion as has been done in [1, 9] cannot be employed for our pinning model. The normalization in the Wick-ordered partition function (1.11) results in a Wick exponential, which allows us to perform a Wick expansion (see Section 3 and also the discussion in [21]). We point out that the Wick-ordered partition function (1.11) reduces to [9, equation (3.3)] when the disorders are i.i.d., noting that $\frac{1}{2}\mathrm{Var}_{\omega}(\sum_{x\in\Omega_{N}}\beta\omega_{Nx}\delta_{x})=\frac{1}{2}\beta^{2}\sum_{n=1}^{N}\mathbf{1}_{\{n\in\tau\}}$. After addressing the Wick-ordered partition function, we turn to the original partition function $Z_{N,\beta}^{\omega,h}$ given in (1.9), for which we carry out a Taylor expansion. The finite-order chaos in the Wick and Taylor expansions is connected by the Hu–Meyer formula (see Eq. (2.10) in Section 2).
@@
[…]
@@
Let $\phi_{k}(t_{1},\dots,t_{k})$ be a symmetric function. For any given $(t_{1},\cdots,t_{k})\in(0,1)^{k}$ with $t_{i}\neq t_{j}$ for any $i\neq j$ (if there exist $i\neq j$, s.t. $t_{i}=t_{j}$, we set $\phi_{k}=0$), there exists a unique permutation $\sigma$ on $\llbracket k\rrbracket$ such that $0<t_{\sigma(1)}<t_{\sigma(2)}<\cdots<t_{\sigma(k)}<1$. We define $\phi_{k}$ as follows:
[…]
where $C_{\alpha}$ is the same as that in (1.4).
Now we present our main results for the Wick-ordered partition function $\tilde{Z}_{N,\beta}^{\omega,h}$ given in (1.11) and for the original partition function $Z_{N,\beta}^{\omega,h}$ given in (1.9), respectively. The weak limits $\tilde{Z}_{\hat{\beta},\hat{h}}$ in (1.14) and $Z_{\hat{\beta},\hat{h}}$ in (1.15) involve multiple stochastic integrals in the sense of Skorohod and Stratonovich (see Section 2 for their definitions), respectively. Roughly speaking, the multiple Skorohod integrals in (1.14) (resp. the multiple Stratonovich integrals in (1.15)) can be defined as the limit of Riemann sums, where the products are taken as Wick products (resp. ordinary products). Although multiple Skorohod and Stratonovich integrals are related through the Hu–Meyer formula (2.10), it is not straightforward to establish a direct connection between $\tilde{Z}_{\hat{\beta},\hat{h}}$ and $Z_{\hat{\beta},\hat{h}}$.
@@
Let the renewal process $\tau$ and the disorder $\omega$ be given as in Section 1.1. Let H be the parameter in (1.6)-(1.7) with $H\in(\frac{1}{2},1)$. Recall the Wick-ordered partition function $\tilde{Z}^{\omega,h}_{N,\beta}$ given in (1.11) and the scalings (1.12). We have the following three cases for different $\alpha>0$ in (1.1).
Case (1): If $\alpha>\frac{1}{2}$, then for any $\hat{\beta}>0$ and $\hat{h}\in\mathbb{R}$, we have that
$\tilde{Z}^{\omega,h_{N}}_{N,\beta_{N}}\overset{(d)}{\longrightarrow}\tilde{Z}_{\hat{\beta},\hat{h}}~,\text{ as }N\to\infty.$
Here $\tilde{Z}_{\hat{\beta},\hat{h}}$ is given by the $L^{2}$-convergent series:
where $\phi_{k}$ is given by (1.13), $\prod\diamond$ is the Wick product, $\dot{W}$ is Gaussian noise with $\mathbb{E}[\dot{W}(t)\dot{W}(s)]=|t-s|^{2H-2}$ and the stochastic integral is of Skorohod type.
Case (2): If $\alpha=\frac{1}{2}$ and we further assume that the slowly varying function $L(\cdot)$ in (1.1) satisfies $0<\inf_{n\in\mathbb{N}}L(n)\leq\sup_{n\in\mathbb{N}}L(n)<+\infty$, then there exist some $\hat{\beta}_{c},\hat{h}_{c}\in(0,+\infty)$, such that for $\hat{\beta}\in(0,\hat{\beta}_{c})$ and $|\hat{h}|<\hat{h}_{c}$, we have the same weak convergence as in Case (1), that is,
@@
where $\tilde{Z}_{\hat{\beta},\hat{h}}$ is given by the $L^{2}$-convergent series (1.14).
Case (3): If $\alpha<\frac{1}{2}$, then for any $\hat{\beta}>0$ and $\hat{h}\in\mathbb{R}$, we have
$\lim\limits_{N\to\infty}\mathbb{E}\Big[\Big(\tilde{Z}^{\omega,h_{N}}_{N,\beta_{N}}\Big)^{2}\Big]=+\infty.$
Remark 1.2.
Note that $\alpha=\tfrac{1}{2}$ is the threshold between the $L^{2}$ regime and the non-$L^{2}$ regime, which will be further discussed in Remark 3.3. The case $\alpha<\tfrac{1}{2}$ will be addressed in Remark 1.6 below.
@@
Let the renewal process $\tau$ and the disorder $\omega$ be given as in Section 1.1. Let H and $\alpha$ be parameters satisfying $H\in(\frac{1}{2},1)$ and $\alpha+H>\frac{3}{2}$. Recalling the partition function $Z^{\omega,h}_{N,\beta}$ given in (1.9) and the scalings (1.12), we have that for any $\hat{\beta}>0$ and $\hat{h}\in\mathbb{R}$,
$Z^{\omega,h_{N}}_{N,\beta_{N}}\overset{(d)}{\longrightarrow}Z_{\hat{\beta},\hat{h}}~,~as~N\to\infty.$
Here $Z_{\hat{\beta},\hat{h}}$ is given by the $L^{2}$-convergent series:
@@
Assuming the conditions $H\in(\frac{1}{2},1)$ and $\alpha+H>\frac{3}{2}$ as in Theorem 1.3, the rescaled original partition function $Z^{\omega,h_{N}}_{N,\beta_{N}}$ is actually uniformly bounded in $L^{1}$ (see Section 4.3), and thus no normalization is needed in this situation.
In contrast, under the weaker conditions $H\in\bigl(\tfrac{1}{2},1\bigr)$ and $\alpha\geq\frac{1}{2}$ assumed in Cases (1) and (2) in Theorem 1.1, Wick normalization is employed to obtain weak convergence to a nontrivial random variable. In our setting with colored Gaussian noise, Wick normalization provides more convenience for computations than the “standard” normalization. Furthermore, since the polymer measure acts as a random measure of the underlying renewal process, it is natural to adopt Wick normalization so that the normalized partition function has a mean equal to the homogeneous partition function (i.e., $Z_{N,\beta=0}^{\omega,h}$), which is $1$ when $h=0$. This viewpoint has also been adopted very recently in [28] for the parabolic Anderson model on $\mathbb{R}^{2}$, where the classical Skorohod integral was extended to the $L^{1}$ setting and the global existence of an $L^{1}$ Skorohod solution of the stochastic heat equation with multiplicative planar white noise was proved.
For the directed polymer model in time-correlated random environment, the continuum Skorohod (resp. Stratonovich) partition function corresponds to the Skorohod (resp. Stratonovich) solution of the stochastic heat equation driven by time-colored Gaussian noise (see [29]). However, the literature on SPDEs related to pinning models in random environments is very limited. A stochastic heat equation with multiplicative white noise, corresponding to the marginally relevant pinning model with $\alpha=\tfrac{1}{2}$ in an independent random environment, was recently studied in [32]. In our forthcoming work [22], we investigate a class of fractional stochastic heat equations in the sense of Skorohod (resp. Stratonovich) associated with the Skorohod (resp. Stratonovich) continuum partition functions of random pinning models.
Remark 1.5.
The results in Theorems 1.1 and 1.3 also hold for the conditioned partition functions
$\tilde{Z}_{N}^{\text{c}}=\tilde{Z}_{N,\beta_{N}}^{\omega,h_{N},\text{c}}:=\mathbb{E}_{\tau}\Big[\exp\Big\{\sum_{x\in\Omega_{N}}(\beta\omega_{Nx}+h)\delta_{x}-\frac{1}{2}\mathrm{Var}\Big[\sum_{x\in\Omega_{N}}(\beta\omega_{Nx})\delta_{x}\Big]\Big\}\Big|N\in\tau\Big],~\text{and}$
@@
with the integrand $\phi_{k}$ in (1.14) and (1.15) being replaced by a symmetric function $\phi_{k}^{c}$ with the same symmetrization as for $\phi_{k}$, which for $0<t_{1}<\cdots<t_{k}<1$ is defined by
[…]